I'm trying to set this problem up
The lenghth of a rectangle is six cm more than one and one-half times the width.The perimeter of the rectangle if fifty-six cm.find the dimensions.
I put l=6+1and1/2
56=2w+2l
I don't think that's right though

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

Second equation is fine.
L=6+(1.5W)
56=2L+2W

Answer by stanbon(75887)

The length of a rectangle is six cm more than one and one-half times the width.The perimeter of the rectangle if fifty-six cm.find the dimensions.
--------
Let width be "W":
Then L = (3/2)W + 6
-----
Perimeter = 2(L + W)
56 = 2((3/2)W+6 + W)
---
28 = (5/2)W + 6
---
(5/2)W = 22
------
Width = (2/5)(22) = 44/5 = 8.8 cm
-----
Lenght = (3/2)(8.8) + 6 = 19.2 cm
====================================
Cheers,
Stan H.
================